# NeuroLifting: Neural Inference on Markov Random Fields at Scale

### v2

@@ Abstract
Inference in large-scale Markov Random Fields (MRFs) remains challenging, with traditional approximate like belief propagation and exact methods such as the Toulbar2 solver often struggling to balance efficiency and solution quality at scale. This paper presents NeuroLifting, a novel approach that uses Graph Neural Networks (GNNs) to reparameterize MRF decision variables, enabling standard gradient descent optimization. By extending lifting techniques through neural networks, NeuroLifting achieves efficient, parallelizable optimization with a smooth loss landscape. Empirical results show NeuroLifting matches Toulbar2’s solution quality on moderate scales while outperforming approximate methods. Notably, on large-scale MRFs, it demonstrates superior solutions compared to baselines with linear computational complexity growth, marking a significant advance in scalable MRF inference.
@@ 1 Introduction
Markov Random Fields stand as a fundamental computational paradigm for modeling complex dependencies among a large collection of variables, permeating a variety of domains such as computer vision [1, 2], natural language processing [3, 4, 5], and network analysis [6, 7]. MRF’s capacity to encode intricate probabilistic interactions underscores its widespread utility. However, unraveling the optimal configurations in high-dimensional settings remains a formidable task owing to the inherent computational complexity involved.
Traditional inference methodologies for MRFs bifurcate into approximate and exact strategies, each with its own set of advantages and limitations. Approximate inference techniques, such as belief propagation [8, 9] and mean field [10, 11] approximations, strive for computational efficiency but often at the expense of solution quality, particularly as the scale of the problem escalates. Conversely, exact inference methods, epitomized by the Toulbar2 solver [12, 13], aspire to optimality but are frequently hampered by exponential time complexities that render them infeasible for large-scale MRFs.
Despite significant advances, achieving a harmonious balance between efficiency and solution quality in large-scale MRF inference has remained a long-standing unmet challenge. This paper addresses this pivotal issue through the introduction of “NeuroLifting” – a neural-network-driven paradigm that extends traditional lifting technique in the context of optimization [14, 15, 16]. NeuroLifting is a novel approach that reimagines MRF inference by leveraging the potency of GNNs alongside gradient-based optimization techniques.
The core innovation of NeuroLifting lies in the reparameterization of the decision variables within MRFs utilizing a randomly initialized GNN. While some recent heuristics succeeded in utilizing GNNs for solving combinatorial problems [17, 18], an effective adaptation to MRF inference remains opaque. Besides, they generally lack an in-depth understanding of how GNNs facilitate downstream computation. In this paper, we for the first time bridge such practice to traditional lifting techniques, and further demonstrate that by harnessing the continuous and smooth loss landscape intrinsic to neural networks, NeuroLifting simplifies the optimization process for large-scale MRFs, enabling enhanced parallelization and performance on GPU devices.
@@
In summary, the contributions of this paper are threefold. 1) Methodical design: we present NeuroLifting as an innovative and practical solution to the enduring challenge of efficient and high-quality inference in large-scale MRFs; 2) Non-parametric lifting: we extend the concept of lifting from traditional optimization practices into a modern neural network framework, thereby offering a fresh lens through which to tackle large-scale inference problems; 3) Significant performance: NeuroLifting achieved significant performance improvement over existing methods, showing remarkable scalability and efficiency in real-world scenarios.
@@ 2 Related work
Lifting techniques have significantly impacted optimization, especially for combinatorial problems and algorithm enhancement [19]. These methods transform problems into higher-dimensional spaces for better representation and solutions. Applications include mixed 0-1 integer programming as shown by [20] and MIP problems with primal cutting-plane algorithms as demonstrated by [21]. Integration with variable upper bound constraints has proven effective for problems like Knapsack [22]. The techniques have expanded to robust optimization [23, 24] and shown promise when combined with Newton’s method for NLPs [14].
Unsupervised GNNs for Combinatorial Optimization.
@@
Graph Neural Networks have demonstrated their power in optimization [25, 26], with recent unsupervised GNN advancements showing effectiveness in combinatorial optimization. Unsupervised GNNs can learn meaningful representations of nodes and edges without labeled data, effectively capturing combinatorial structure as shown by [27]. This approach has proven particularly valuable for problems like the Traveling Salesman Problem [28, 29], Vehicle Routing Problem [30], and Boolean satisfiability problem [17]. Efficient solutions for Max Independent Set and Max Cut problems were also demonstrated by [18]. However, the loss functions may lack flexibility in handling higher-order relationships beyond edges.
@@
MRF and Inference. The maximum a posterior (MAP) problem of MRFs is finding the best configuration that could minimize the energy function which is a NP-hard problem. Currently, the popular methods include variants of belief propagation [31, 32, 33] and tree-reweighted message passing (TRBP) [9, 34] and a generalization of TRBP which is suitable for high-order MRFs(SRMP) [35].
@@
Neural Networks with supervised learning remain the mainstream approach for MRF problems. GNNs were used for MRF inference by [36], outperforming LBP and TRBP but limited to 16 nodes. BP information was integrated into GNNs to formulate a neural factor graph by [37], surpassing traditional BP on graphs under 100 nodes. GNN message passing rules were modified by [38] to align with BP properties, showing better performance than LBP on graphs up to 196 nodes. A neural version of the Max-product algorithm was proposed by [39]. Variational MPNN for MAP problems on 9-13 node graphs was introduced by [40]. Researchers also explored CRFs. The semi-supervised method by [41] used CRF to enhance GNN classification, further implemented by [42]. However, obtaining optimal configurations in large-scale MRFs remains challenging. Unsupervised and self-supervised learning offer alternative approaches. Learning to optimize principles [43] were applied to MRF problems, while the Augmented Lagrangian Method was employed by [44] as loss function for self-supervised models solving CMPE problems. A self-supervised approach for binary node MMAP problems was proposed by [45]. These methods require specially designed loss functions and are all limited to smaller instances.
@@ 3 Preliminary
$\mathbb{P}(x_{i}|X\backslash\{x_{i}\})=\mathbb{P}(x_{i}|\{x_{j}\}~{}~{}\text{for}~{}~{}i,j\in C_{k}~{}~{}\text{for}~{}~{}C_{k}\in\mathcal{C}).$
In this paper, we consider the MAP estimate of Eq. 1, which requests optimizing Eq. 1 via $X^{*}=\min_{X}E(X)$. One can consult [46] for more details.
Graph Neural Networks.  GNNs represent a distinct class of neural network architectures specifically engineered to process graph-structured data [47, 48, 49, 50]. In general, when addressing a problem involving a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{E}$ is the edge set, GNNs utilize both the graph $\mathcal{G}$ and the initial node representations $\{h_{i}^{(0)}\in\mathbb{R}^{d}|\forall i\in\mathcal{V}\}$ as inputs, where $d$ is the dimension of initial features. Assuming the total number of GNN layers to be $K$, at the $k$-th layer the graph convolutions typically read:
@@
Optimization with Lifting.
Lifting is a sophisticated technique employed in the field of optimization to address and solve complex problems by transforming them into higher-dimensional spaces [51, 52]. By introducing auxiliary variables or constraints, lifting serves to reformulate an original optimization problem into a more tractable or elucidated form, often making the exploration of optimal solutions more accessible. In the context of MRFs, lifting can be utilized to transform inference problems into higher dimensions where certain properties or symmetries associated with specific MRF problems are more easily exploitable  [9, 53, 16]. However, a principled lifting technique is still lacking for generalized MRFs.
@@ 4.2 Prepossessing
Vectorizing the energy function. The transformed energy function $E(X)$ will serve as the loss function guiding the training of the neural networks. In Section. 4.4, we will detail the transformation process and discuss how to effectively utilize it. Note the values of these functions can be pre-evaluated and repeatedly used during the training process. Therefore, we employ a look-up table to memorize all function values with discrete inputs. For unary energies, we denote the vectorized unary energy of variable $x_{i}$ as $\phi(x_{i})$, where the $n$-th element corresponds to $\theta_{i}(x_{i}=n)$. Similarly, we represent the clique energy for clique $C_{k}$ using the tensor $\psi(\{x_{l}|\forall x_{l}\in C_{k}\})$. This tensor can be derived using the same conceptual framework; for instance, the element $\psi(x_{i},x_{j},x_{k})$ at position $(0,2,4)$ corresponds to the value of $\theta_{\{i,j,k\}}(x_{i}=0,x_{j}=2,x_{k}=4)$
Padding node embeddings & energy terms and Masking.
GNNs typically require all node embeddings to be of the same dimension, meaning that the embeddings $h^{(K)}$ at $K$-th layer must share the same size. However, in general MRFs, the variables often exhibit different numbers of states. While traditional belief-propagation-based methods can easily manage such variability, adapting GNNs to handle these discrepancies is less straightforward.
To address this mismatch, we employ padding strategy – a common technique used to handle varying data lengths. This strategy is applied to both the node embeddings and the unary and pairwise (or clique) energies, to ensure consistent embedding dimensions. Concretely, we assign virtual states to the nodes whose state number is less than $|\mathcal{X}|$. Then, we assign energies to those padded labels with the largest value of the original energy term. The schematic diagram of the padding procedure could be found in Appendix. A. This approach of assigning high energies to the padded labels serves to discourage the model from selecting these padded states, thereby incentivizing it to choose the original, non-padded states with lower energies. We employ a masking strategy to exclude padded regions(using -inf as the mask), thereby ensuring these artificially added areas neither participate in the selection process nor significantly affect the loss computation.
@@
Although utilizing uniform large values(e.g., padding with inf) for energy padding is theoretically viable, this approach introduces significant computational bias. Specifically, this padding methodology causes substantial deviation between the training loss and the true energy metrics, as minimal variations in padded regions disproportionately influence the loss function. Such distortion impedes effective monitoring of the training dynamics. As empirically validated in Section. 5, our proposed padding scheme demonstrates superior performance by maintaining a high degree of consistency between the training loss and the true energy measurements, thereby ensuring more reliable model optimization.
@@ 4.3 GNNs as Non-parametric Lifting
In this section, we detail how NeuroLifting generates features that capture the hidden information of the given MRF and solves the original MAP problem by optimizing in a high-dimensional parameter space. As mentioned in Section. 4.2, we initially generate learnable feature vectors randomly using an encoder that embeds all nodes, transforming the integer decision variables into $d_{l}$-dimension vectors $h_{i}^{(0)}\in\mathbb{R}^{d_{l}}$ for node $i$, where $d_{l}$ is a hyper-parameter representing the dimension after lifting.
The intuition for utilizing GNNs in the implementation of lifting techniques is inspired by LBP [31]. When applying LBP for inference on MRFs, the incoming message $M_{ji}$ to node $i$ from node $j$ is propagated along the edges connecting them. Node $i$ can then update its marginal distributions according to the formula in Eq. 4 where $exp(-\phi(x_{i}))$ is the unary potential function.
$p^{\textbf{posterior}}(x_{i}|X\backslash\{x_{i}\})=\exp(-\phi(x_{i}))\prod\limits_{(i,j)\in\mathcal{E}}\sum\limits_{x_{j}}M_{ji}$
@@
Graph convolutions should intuitively treat adjacent nodes equally, consistent with the principle in MRFs, where the information collected from neighbors is processed equally. Typical GNNs are summarized in Table. 8 from Appendix. 8, where $\operatorname{deg}(i)$ is the degree of node $i$, $\alpha_{i,j}$ is the attention coefficients, and $|\mathcal{N}(i)|$ is the neighborhood size of node $i$. According to the influence of neighbors, they can be classified into three categories: 1) neighborhood aggregation with normalizations (e.g., GCN [47] normalize the influence by node degrees), 2) neighborhood aggregation with directional biases (e.g., GAT [50] learn to select the important neighbors via an attention mechanism), and 3) neighborhood aggregation without bias (e.g., GraphSAGE [48] directly aggregate neighborhood messages with the same weight). Therefore, we select the aggregator in GraphSAGE as our backbone for graph convolutions. The performance of these GNN backbones on our MRF datasets is shown in Fig. 5 in Appendix. G.
Another primary characteristic of MRFs is its ability to facilitate information propagation across the graph through local connections. This means that even though the interactions are defined locally between neighboring nodes, the influence of a node can extend far beyond its immediate vicinity. As a result, MRFs can effectively capture global structure and dependencies within the data. We thus use Jumping Knowledge [54] to leverage different neighborhood ranges. By doing so, features representing local properties can utilize information from nearby neighbors, while those indicating global states may benefit from features derived from higher layers.

### v3

@@ Abstract
Inference in large-scale Markov Random Fields (MRFs) remains challenging, with traditional approximate methods like belief propagation and exact methods such as the Toulbar2 solver often struggling to balance efficiency and solution quality at scale. This paper presents NeuroLifting, a novel approach that uses Graph Neural Networks (GNNs) to reparameterize MRF decision variables, enabling standard gradient descent optimization. By extending lifting techniques through neural networks, NeuroLifting achieves efficient, parallelizable optimization with a smooth loss landscape. Theoretically, we provide a rigorous analysis proving that our neural relaxation is consistent with the discrete objective and constructs a connected parameter space that allows the optimization to bypass high-energy barriers. Empirical results show NeuroLifting matches Toulbar2’s solution quality on moderate scales while outperforming approximate methods. Notably, on large-scale MRFs, it demonstrates superior solutions compared to baselines with linear computational complexity growth, marking a significant advance in scalable MRF inference.
Machine Learning, ICML
@@ 1 Introduction
Markov Random Fields stand as a fundamental computational paradigm for modeling complex dependencies among a large collection of variables, permeating a variety of domains such as computer vision (Wang et al., 2013; Su et al., 2021), natural language processing (Almutiri & Nadeem, 2022; Ammar et al., 2014; Lin et al., 2020), and network analysis (Wu et al., 2020; Yunfei Ma & Razavi, 2022). MRF’s capacity to encode intricate probabilistic interactions underscores its widespread utility. However, unraveling the optimal configurations in high-dimensional and large-scale settings remains a formidable task owing to the inherent computational complexity involved.
Traditional inference methodologies for MRFs bifurcate into approximate and exact strategies, each with its own set of advantages and limitations. Approximate inference techniques, such as belief propagation (Pearl, 2022; Wainwright et al., 2005) and mean field (Saito et al., 2012; Zhang, 1993) approximations, strive for computational efficiency but often at the expense of solution quality, particularly as the scale of the problem escalates. Conversely, exact inference methods, epitomized by the Toulbar2 solver (De Givry, 2023; Hurley et al., 2016), aspire to optimality but are frequently hampered by exponential time complexities that render them infeasible for large-scale MRFs.
Despite significant advances, achieving a harmonious balance between efficiency and solution quality in large-scale MRF inference has remained a long-standing unmet challenge. This paper addresses this pivotal issue through the introduction of “NeuroLifting” – a neural-network-driven paradigm that extends traditional lifting technique in the context of optimization (Albersmeyer & Diehl, 2010; Balas & Perregaard, 2002; Bauermeister et al., 2022). NeuroLifting is a novel approach that reimagines MRF inference by leveraging the potency of GNNs alongside gradient-based optimization techniques.
The core innovation of NeuroLifting lies in the reparameterization of the decision variables within MRFs utilizing a randomly initialized GNN. While some recent heuristics succeeded in utilizing GNNs for solving combinatorial problems (Cappart et al., 2023; Schuetz et al., 2022), an effective adaptation to MRF inference remains opaque. Besides, they generally lack an in-depth understanding of how GNNs facilitate downstream computation. In this paper, we for the first time bridge such practice to traditional lifting techniques, and further theoretically demonstrate that by harnessing the continuous and smooth loss landscape intrinsic to neural networks, NeuroLifting simplifies the optimization process for large-scale MRFs, enabling enhanced parallelization and performance on GPU devices.
@@
In summary, the contributions of this paper are threefold. 1) Methodical design: we present NeuroLifting, which extends the concept of lifting from traditional optimization practices into a modern neural network framework, as an innovative and practical solution to the enduring challenge of efficient and high-quality inference in large-scale MRFs; 2) Theoretical Justification: we provide rigorous analysis proving that our neural relaxation is consistent with the discrete objective, and demonstrate that the over-parameterized space constructs low-energy tunnels allowing optimization to bypass high-energy barriers between local minima; 3) Significant performance: NeuroLifting achieves significant performance improvement over existing methods, showing remarkable scalability and efficiency in real-world scenarios.
@@ 2 Related work
Lifting in Optimization. Lifting techniques have significantly impacted optimization, especially for combinatorial problems and algorithm enhancement (Marchand et al., 2002). These methods transform problems into higher-dimensional spaces for better representation and solutions. Applications include mixed 0-1 integer programming as shown by Balas et al. (1993) and MIP problems with primal cutting-plane algorithms as demonstrated by Dey & Richard (2008). Integration with variable upper bound constraints has proven effective for problems like Knapsack (Shebalov et al., 2015). The techniques have expanded to robust optimization (Georghiou et al., 2020; Bertsimas et al., 2019) and shown promise when combined with Newton’s method for NLPs (Albersmeyer & Diehl, 2010).
@@
Unsupervised GNNs for Combinatorial Optimization. Graph Neural Networks have demonstrated their power in optimization (Yu et al., 2019; Ying et al., 2024), with recent unsupervised GNN advancements showing effectiveness in combinatorial optimization. Unsupervised GNNs can learn meaningful representations of nodes and edges without labeled data, effectively capturing combinatorial structure as shown by Peng et al. (2021). This approach has proven particularly valuable for problems like the Traveling Salesman Problem (Gaile et al., 2022; Min et al., 2023), Vehicle Routing Problem (Wu et al., 2024), and Boolean satisfiability problem (Cappart et al., 2023). Efficient solutions for Max Independent Set and Max Cut problems were also demonstrated by Schuetz et al. (2022). However, the loss functions may lack flexibility in handling higher-order relationships beyond edges.
@@
MRF and Inference. The maximum a posterior (MAP) problem of MRFs is finding the best configuration that could minimize the energy function which is a NP-hard problem. Currently, the popular methods include variants of belief propagation (Weiss & Freeman, 2001; Felzenszwalb & Huttenlocher, 2004; Frey & Mackay, 2002) and tree-reweighted message passing (TRBP) (Wainwright et al., 2005; Kolmogorov, 2006) and a generalization of TRBP which is suitable for high-order MRFs(SRMP) (Kolmogorov, 2015).
@@
Learning for MRFs. Neural Networks with supervised learning remain the mainstream approach for MRF problems. GNNs were used for MRF inference by Yoon et al. (2019), outperforming LBP and TRBP but limited to 16 nodes. BP information was integrated into GNNs to formulate a neural factor graph by Garcia Satorras & Welling (2021), surpassing traditional BP on graphs under 100 nodes. GNN message passing rules were modified by Kuck et al. (2020) to align with BP properties, showing better performance than LBP on graphs up to 196 nodes. A neural version of the Max-product algorithm was proposed by Zhang et al. (2020). Variational MPNN for MAP problems on 9-13 node graphs was introduced by Cui et al. (2022). Researchers also explored CRFs. The semi-supervised method by Qu et al. (2019) used CRF to enhance GNN classification, further implemented by Tang et al. (2021). However, obtaining optimal configurations in large-scale MRFs remains challenging. Unsupervised and self-supervised learning offer alternative approaches. Learning to optimize principles (Nair et al., 2021) were applied to MRF problems, while the Augmented Lagrangian Method was employed by Arya et al. (2024b) as loss function for self-supervised models solving CMPE problems. A self-supervised approach for binary node MMAP problems was proposed by Arya et al. (2024a). These methods require specially designed loss functions and are all limited to smaller instances.
@@ 3 Preliminary
$\begin{aligned} \mathbb{P}(X)=\frac{1}{Z}\exp(-E(X))=\frac{1}{Z}\exp\left(-\sum_{i\in\mathcal{V}}\theta_{i}(x_{i})-\sum_{C_{k}\in\mathcal{C}}\theta_{C_{k}}(\{x_{l}|\forall x_{l}\in C_{k}\})\right)\end{aligned}$
@@
In this paper, we consider the MAP estimate of Eq. 1, which requests optimizing Eq. 1 via $X^{*}=\min_{X}E(X)$. One can consult Koller & Friedman (2009) for more details.
Graph Neural Networks.  GNNs represent a distinct class of neural network architectures specifically engineered to process graph-structured data (Kipf & Welling, 2017; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2018). In general, when addressing a problem involving a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{E}$ is the edge set, GNNs utilize both the graph $\mathcal{G}$ and the initial node representations $\{h_{i}^{(0)}\in\mathbb{R}^{d}|\forall i\in\mathcal{V}\}$ as inputs, where $d$ is the dimension of initial features. Assuming the total number of GNN layers to be $K$, at the $k$-th layer the graph convolutions typically read:
@@
Optimization with Lifting.  Lifting is a sophisticated technique employed in the field of optimization to address and solve complex problems by transforming them into higher-dimensional spaces (Balas, 2005; Papadimitriou & Steiglitz, 1982). By introducing auxiliary variables or constraints, lifting serves to reformulate an original optimization problem into a more tractable or elucidated form, often making the exploration of optimal solutions more accessible. In the context of MRFs, lifting can be utilized to transform inference problems into higher dimensions where certain properties or symmetries associated with specific MRF problems are more easily exploitable  (Wainwright et al., 2005; Globerson & Jaakkola, 2007; Bauermeister et al., 2022). However, a principled lifting technique is still lacking for generalized MRFs.
@@ 4.2 Prepossessing
Vectorizing the energy function. The transformed energy function $E(X)$ will serve as the loss function guiding the training of the neural networks. In Section. 4.4, we will detail the transformation process and discuss how to effectively utilize it. Note the values of these functions can be pre-evaluated and repeatedly used during the training process. Therefore, we employ a look-up table to memorize all function values with discrete inputs. For unary energies, we denote the vectorized unary energy of variable $x_{i}$ as $\phi(x_{i})$, where the $n$-th element corresponds to $\theta_{i}(x_{i}=n)$. Similarly, we represent the clique energy for clique $C_{k}$ using the tensor $\psi(\{x_{l}|\forall x_{l}\in C_{k}\})$. This tensor can be derived using the same conceptual framework; for instance, the element $\psi(x_{i},x_{j},x_{k})$ at position $(0,2,4)$ corresponds to the value of $\theta_{\{i,j,k\}}(x_{i}=0,x_{j}=2,x_{k}=4)$ .
Padding node embeddings & energy terms and Masking. GNNs typically require all node embeddings to be of the same dimension, meaning that the embeddings $h^{(K)}$ at $K$-th layer must share the same size. However, in general MRFs, the variables often exhibit different numbers of states. While traditional belief-propagation-based methods can easily manage such variability, adapting GNNs to handle these discrepancies is less straightforward.
To address this mismatch, we employ padding strategy – a common technique used to handle varying data lengths. This strategy is applied to both the node embeddings and the unary and pairwise (or clique) energies, to ensure consistent embedding dimensions. Concretely, we assign virtual states to the nodes whose state number is less than $|\mathcal{X}|$. Then, we assign energies to those padded labels with the largest value of the original energy term. The schematic diagram of the padding procedure could be found in Appendix A. This approach of assigning high energies to the padded labels serves to discourage the model from selecting these padded states, thereby incentivizing it to choose the original, non-padded states with lower energies. We employ a masking strategy to exclude padded regions(using -inf as the mask), thereby ensuring these artificially added areas neither participate in the selection process nor significantly affect the loss computation.
@@
Although utilizing uniform large values(e.g., padding with inf) for energy padding is theoretically viable, this approach introduces significant computational bias. Specifically, this padding methodology causes substantial deviation between the training loss and the true energy metrics, as minimal variations in padded regions disproportionately influence the loss function. Such distortion impedes effective monitoring of the training dynamics. As empirically validated in Section 5, our proposed padding scheme demonstrates superior performance by maintaining a high degree of consistency between the training loss and the true energy measurements, thereby ensuring more reliable model optimization.
@@ 4.3 GNNs as Non-parametric Lifting
In this section, we detail how NeuroLifting generates features that capture the hidden information of the given MRF and solves the original MAP problem by optimizing in a high-dimensional parameter space. As mentioned in Section 4.2, we initially generate learnable feature vectors randomly using an encoder that embeds all nodes, transforming the integer decision variables into $d_{l}$-dimension vectors $h_{i}^{(0)}\in\mathbb{R}^{d_{l}}$ for node $i$, where $d_{l}$ is a hyper-parameter representing the dimension after lifting.
The intuition for utilizing GNNs in the implementation of lifting techniques is inspired by LBP. When applying LBP for inference on MRFs, the incoming message $M_{ji}$ to node $i$ from node $j$ is propagated along the edges connecting them. Node $i$ can then update its marginal distributions according to the formula in Eq. 4 where $exp(-\phi(x_{i}))$ is the unary potential function.
[…]
@@
Graph convolutions should intuitively treat adjacent nodes equally, consistent with the principle in MRFs, where the information collected from neighbors is processed equally. Typical GNNs are summarized in Table 8 from Appendix 8, where $\operatorname{deg}(i)$ is the degree of node $i$, $\alpha_{i,j}$ is the attention coefficients, and $|\mathcal{N}(i)|$ is the neighborhood size of node $i$. According to the influence of neighbors, they can be classified into three categories: 1) neighborhood aggregation with normalizations (e.g., GCN (Kipf & Welling, 2017) normalize the influence by node degrees), 2) neighborhood aggregation with directional biases (e.g., GAT (Veličković et al., 2018) learn to select the important neighbors via an attention mechanism), and 3) neighborhood aggregation without bias (e.g., GraphSAGE (Hamilton et al., 2017) directly aggregate neighborhood messages with the same weight). Therefore, we select the aggregator in GraphSAGE as our backbone for graph convolutions. The performance of these GNN backbones on our MRF datasets is shown in Fig. 5 in Appendix H.
Another primary characteristic of MRFs is its ability to facilitate information propagation across the graph through local connections. This means that even though the interactions are defined locally between neighboring nodes, the influence of a node can extend far beyond its immediate vicinity. As a result, MRFs can effectively capture global structure and dependencies within the data. We thus use Jumping Knowledge (Xu et al., 2018) to leverage different neighborhood ranges. By doing so, features representing local properties can utilize information from nearby neighbors, while those indicating global states may benefit from features derived from higher layers.